OpenStudy (anonymous):

Give the slope and y intercept of each line and graph it. y=3x-1 slope: 3 y: -1 But how would I go about getting the points for graphing?

OpenStudy (anonymous):

a simple solution for any type of graph is is to choose a number for x and solve for y for example if x = 0 then y=3(0)-1 which is -1 so x one set of points is x=1,y=-1

OpenStudy (anonymous):

i mean x=0 y=-1

OpenStudy (anonymous):

?

OpenStudy (anonymous):

you can substitute any number for x and figure out what y is. if Y=3x-1 you can pick any number for x and solve for y. for example if you pick 1 for x you are left with the equation y=3(1)-1 notice I put 1 in the place of x then you use algebra to solve for Y . y=3(1)-1 y=2 then the point on the graph is the value you picked for x which is 1 and the solution for y which is 2 the set of points is (1,2)